-0.5(8 - 4t) = 2(2 + t)
Answer by mananth(16946)

-0.5(8 - 4t) = 2(2 + t)
..
-4+2t= 4+2t
-2t
-4+2t-2t=2t-2t+4
-4=4
...
no solution
...